# On Modeling and Solving the Boltzmann Equation

### v1

@@ Abstract
Keywords: Boltzmann Equation, Transport Equation, Radiative Transfer Equation, Analytical Discrete Ordinates (ADO) Method
@@ 1 Introduction
The Boltzmann equation (BE) (Boltzmann, 1872) has been the reason and motivation for relevant mathematical studies and research over the years (Cercignani, 1988; Villani and Desvillettes, 2005). Its challenging theoretical complexity, combined with a wide variety of current problems that require numerical simulations based on this model, justifies such interest. The fundamental model was derived to describe the particle distribution. In the context of the kinetic theory of gases, this concerns a nonlinear integrodifferential operator. The treatment of the collision integral operator is known to be a significant challenge. For simulation purposes, the Linearized Boltzmann equation (LBE) or kinetic (linear) models derived from the original equation are commonly used. On the other hand, when interactions between particles are neglected and the goal is to describe the transport of a particle in a material medium, the balance of the particles results in a linear model. The linear version of the Boltzmann equation is commonly referred to as the Transport Equation (TE) or Radiative Transfer Equation (RTE), for photon applications.
@@
In this work, we report studies and advances related to the solution of the linear Boltzmann equation in one and two spatial dimensions. We discuss relevant aspects of neutron and photon transport modeling such as in nuclear safeguards, nuclear reactor shielding problems, and optical tomography. In addition, we include a brief report on kinetic models of the rarefied gas dynamics, which are relevant, for instance, in the studies of micro-electro-mechanical systems, and their connection with the linearized Boltzmann equation (LBE).
@@
The already commented complexity of the Boltzmann models reinforces the development of deterministic methods that can provide high-quality weak form solutions. The discretization of variables is necessary in both purely numerical approaches and analytical techniques. In this context, the discrete ordinates method is well-known and is used in established codes (Lewis and Miller, 1984) to deal with angular variables. The solution for selected discrete directions of the particles, along with the numerical approximation of the integral operator by a quadrature scheme, transform the original problem into a system of first-order partial or ordinary differential equations.
@@
In the literature, the discrete ordinates method is often referred to as the $S_{N}$ method (Davison, 1957). In fact, in this case, in addition to the angular discretization, spatial variables are also usually discretized. The work of Chandrasekhar (1960) in the field of astrophysics was a pioneering effort in proposing analytical solutions for the discrete ordinates approximation of the RTE equation in one-dimensional geometry, in terms of the spatial variable. In more recent years, the Analytical Discrete Ordinates (ADO) Method, (Barichello and Siewert, 1999b, 2001), a version of Chandrasekhar’s solution,  was proposed. The technique comprises a simultaneous set of features that made Chandrasekhar’s original ideas computationally viable, as well as expanded the applicability of the method to problems in other areas of transport theory, such as neutron transport and rarefied gas dynamics. It is worth mentioning the features: the use of an arbitrary numerical quadrature scheme defined in the half-range, not restricted to the Gauss-Legendre quadrature scheme; the separation constants are obtained through an eigenvalue problem instead of searching for roots of polynomials; the resulting eigenvalue problem is of half the order of the number of directions; finally, a scaling introduced to deal with the exponential terms in the solution prevents ”overflow”.
@@
For problems in two-dimensional Cartesian geometry, the discrete ordinates approximation reduces the integrodifferential equation to a system of first-order partial differential equations. Numerical integration on the unit sphere becomes an issue. In the case of spatial variables, the so-called nodal methods (Badruzaman, 1985) reduce the complexity of the model, following the idea of finite volume schemes, and are known to have better performance in coarser meshes. The connection between the spatial and angular grids is usually made in the literature through sweep schemes (Lewis and Miller, 1984) that demand high computational time.
@@
In this framework, the ADO formulation, along with nodal schemes (ADO-Nodal) method, was proposed for solving the discrete ordinates approximation of the two-dimensional linear Boltzmann (transport) equation (Barichello et al., 2011; Picoloto et al., 2015; Barichello et al., 2017). The discrete ordinates equations are transversally integrated over a region (node) of the domain, yielding one-dimensional equations for average angular fluxes or intensities, in $x$ and $y$-directions. The ADO method (Barichello and Siewert, 1999b) is then applied to the one-dimensional equations, with approximations for the unknowns (transverse leakage terms) on the contours of the regions, to derive explicit solutions for the spatial variables.
@@
Due to the analytical characteristic of the ADO-Nodal method, as first proposed by Barichello et al. (2011), the integration procedure to generate the nodal equations was defined for the whole interval of the $x$ and $y$ variables definition, respectively. No division of the domain into nodes was carried out. Later, the idea of local solutions was considered for the treatment of heterogeneous media problems (Barichello et al., 2017) since different materials constitute the physical domain.
@@
Among the good features of this formulation, which is a spectral technique, one may cite: the solution of the transverse integrated one-dimensional equations, either in $x$ and $y$-directions are explicitly written in terms of the spatial variables along with the fact that the associated eigenvalue problem is of reduced order, to only half of the number of discrete directions. Significant advantages in comparison with nodal schemes available by then. With these features, the methodology applied to a broad class of neutron and radiation transport problems has advantages in providing fast and accurate solutions. Enhanced performance has been noted in coarser meshes when compared with analogous methodologies.
@@
The scheme does not use sweeping. Linear systems must be solved to establish the general solutions fully. A significant number of discrete directions in the simulations is theoretically recognized as an error control requirement between the discrete ordinate solution and the exact solution (Madsen, 1971). Furthermore, from a physical perspective, problems in highly anisotropic scattering media, for example, also impose such a condition, which means a choice of high-order quadrature schemes to approximate the integral term of the equation. The quadrature order is a factor that directly impacts the linear system. We implemented and analyzed different approaches for solving the linear system associated with the ADO-Nodal formulation. We analyzed the performance of the schemes applied to problems in different areas of particle transport, such as neutrons and radiation, as well as the influence of using different schemes for representing discrete ordinates (Barichello and da Cunha, 2023).
@@
In this work, we present a general overview of the ADO methodology, aiming to highlight advances in the area and contributions to this topic, as already mentioned, of such complexity. We highlight, in particular, different aspects and mathematical tools necessary for the derivation and implementation of the formulation, as well as in the analysis of errors.
@@
The manuscript is organized as follows: the general integrodifferential formulation is introduced in Section 2, for photons (RTE). In Sections 3 and 4, the one-dimensional simpler models are discussed and the ADO formulation is derived. Two-dimensional geometry problems are presented in Section 5, where the multidimensional quadrature schemes are commented and the ADO-Nodal formulation is developed. Next, in Section 6, neutron transport is discussed, highlighting the error in the spatial discretization of the two-dimensional transport equation. Problems in the general area of rarefied gas dynamics, its connection with the LBE, and the application of the ADO method in this area are analyzed in Section 7. A short report on Inverse problems (Section 8), the field, and the final remarks conclude the text.
2 The Integrodifferential Formulation
We begin with the time-independent (RTE), homogeneous media, which results from a balance in the phase space of gains and losses.
@@ 2 The Integrodifferential Formulation
Here $I({\bf r},{\bm{\Omega}})$ is the intensity of the radiation in ${\bf r}=(x,y,z)$ along the direction $\bm{\Omega}$}; ${\bm{\Omega}}=(\mu,\eta,\xi)$ represents the direction of the particle as a vector in the unit sphere $S$, with $\mu^{2}+\eta^{2}+\xi^{2}=1$, and (see Fig (1))
@@
where $\Theta$, the phase-angle, is the angle between incident and scattered directions; $C_{l}$ are the expansion coefficients; $L$ refer to as the degree of anisotropy of the medium; $P_{l}$ are the Legendre polynomial of lthl-th degree. Using the Addition theorem for the Legendre Polynomials (Stegun, 1965)
@@
The Henyey-Greenstein function is widely used to model scattering in biological tissues, as in optical tomography applications. In this case, $C_{l}=(2l+1)g^{l}$, being $g$ the assimetry factor, $g\in(-1,1)$. That way $g=0$ indicates isotropic scattering while $0<g<1$ indicates forward scattering and 1<g<0-1<g<0: indicates backward scattering. On the other hand, the exact representation of the Henyey-Greenstein function is of the form
@@
To complete the formulation, Eq. (1) is supplemented with boundary conditions. In this work, we introduce them later on in the text.
3 The “Simplest” Problem
For an introduction to transport theory, it may be helpful to start with the simplest possible geometry: a one-dimensional homogeneous medium with isotropic scattering. The general coordinate system described in Fig. (1) is now restricted to what we see in Fig. (2). Under such conditions we consider the problem given by
@@ 3 The “Simplest” Problem
Here $\tau\in\left(\tau_{a},\tau_{b}\right)$, $\tau$ is the optical thickness, $\mu\in\left[-1,1\right]$ is the cosine of the polar angle as measured from the positive $\tau$-axis and $\varpi$ is the albedo for single scattering. We assume that the incoming intensities, $F_{1}(\mu)$ and $F_{2}(\mu)$ are known on the boundaries. Case’s work (Case, 1960) was a landmark in transport theory, presenting a rigorous mathematical development for addressing such a problem. This procedure is now called Case’s method or the singular eigenfunction expansion method. Known as the only exact solution to this problem, despite some additional results (Case and Zweifel, 1967) in the 60’s and 70’s mainly, it is only viable in idealized cases. However, this solution has been the basis for the development of numerical and semi-analytical methods. More importantly, it emphasizes the difficulty of solving problems of this type, characterized by the presence of singular eigenfunctions, associated to continuous and discrete spectrum, and the need for solutions of singular integral equations in assessing orthogonality conditions.
In other words, to address the fundamental problems in various areas of nuclear science, radiative transfer, and rarefied gas dynamics, alternative approaches needed to be developed. Two approaches stand out: probabilistic, such as the Monte Carlo method, and deterministic, which includes the Discrete Ordinates Method.
In the next section, we present the development of the ADO formulation for a still basic problem, with the same geometry as the previous case, which introduces complexities such as the treatment of anisotropic media and boundary conditions that include reflection. The idea is to show how much simpler (and powerful at the same time) such an approach in comparison with Case’s solution.
@@ 4 The ADO Solution for One-dimensional Geometry Problems
In this work, we apply the Analytical Discrete Ordinates (ADO) Method (Barichello and Siewert, 1999b) to solve the slab problem defined by Eqs.(7) to (9). Since the ADO formulation was derived in detail for similar models in earlier papers (Barichello and Siewert, 2001; Barichello, 2011; Groff et al., 2021), we only briefly develop the main steps relevant here. We begin with the equation valid for anisotropic scattering media,
@@
for $\tau\in\left(\tau_{a},\tau_{b}\right)$, $\tau$ is the optical thickness, $\mu\in\left[-1,1\right]$ is the cosine of the polar angle as measured from the positive $\tau$-axis, $\beta_{l}$ are the coefficients in the $L$th-order expansion of the scattering law and $Q(\tau,\mu)$ is an inhomogeneous source term. We assume boundary conditions defined as
@@
for $\mu\in\left(0,1\right]$, which allow for specular and diffuse reflection, with specular and diffuse reflection coefficients, respectively, $\rho_{i}^{s}$ and $\rho_{i}^{d}$, $i=1,2$. We also assume that there are known incident distributions at the boundaries $\tau=\tau_{a}$ and $\tau=\tau_{b}$, given by $F_{1}\left(\mu\right)$ and $F_{2}\left(\mu\right)$, respectively.
We will express the general solution of Eq. (10) as the sum of the homogeneous plus particular solutions. Following Barichello (2011), for developing the homogeneous solution we assume $Q(\tau,\mu)=0$ and rewrite the integral term in equation (10) as
@@
$\pm\mu_{i}\frac{d}{d\tau}\Phi\left(\tau,\pm\mu_{i}\right)+\Phi\left(\tau,\pm\mu_{i}\right)=\frac{\varpi}{2}\sum_{l=0}^{L}\beta_{l}P_{l}\left(\mu_{i}\right)\sum_{k=1}^{N}w_{k}P_{l}\left(\mu_{k}\right)\left[\Phi\left(\tau,\pm\mu_{k}\right)+\left(-1\right)^{l}\Phi\left(\tau,\mp\mu_{k}\right)\right],$ (14)
@@
Equations (14) represent a first-order ordinary differential system of equations. Thus, we seek for exponential solutions of the form
@@
Substituing (15) into equations (14), for $i=1,2,\dots,N$, we obtain
@@
and we add the two set of equations given in (17), to find that
@@
Since $\mu_{i}>0$ for $i=1,2,\dots,N$, we note that the inverse of the matrix $\mathbf{M}$, denoted by $\mathbf{M}^{-1}$, is well defined. Analogously, we subtract the two set of equations in (17), to obtain
@@
We note that the separation constants, $\nu_{j}$, will appear in pairs $\left(\pm\right)$. In addition, the eigenvalue problems (34) and (35) are of half-order in comparison with the ones obtained in the standard discrete ordinates methods, which are based on full-range quadrature schemes. Continuing, from equation (32) we have
@@
from the solution of the eigenvalue problem, along with the elementary solutions, (39) and (40), we can write the discrete ordinates solution for the homogeneous version of equation (10) in the vector form
@@
where $A_{j}$ and $B_{j}$ are the arbitrary coefficientes to be determined next. We note that, in writing the solution in this way, we avoid the computational issue of exponential terms which can lead to overflow (Barichello and Siewert, 2001; Barichello, 2011).
To complete our solution, concerning to particular solutions we follow Barichello et al. (2000). We omit here the details of the derivation which was based on the infinite-medium Green’s function. We express the components of the $N\times 1$ vector
@@
At this point we can represent, in a vector form, the general solution of Eq. (10),
@@
To completely establish the solution we need to determine the superposition coefficients, $A_{j}$ and $B_{j}$. These constants can be calculated as the solution of a $2N\times 2N$ linear system of equations, given below, obtained from the application of the boundary conditions, Eqs. (11) and (12),

### v2

@@ Abstract
keywords:
Boltzmann Equation, Mathematical Physics, Numerical Analysis, Transport Theory, Deterministic Methods.
[…]
Acknowledgements.
The author thanks to CNPq of Brazil for partial financial support to this and other works. The author also wants to express her gratitude to all the co-authors referenced in this work, whose collaboration was fundamental to making it possible to develop the ADO method. Very special thanks to C. E. Siewert.
@@ 1 Introduction
The Boltzmann equation (BE) [24] has been the reason and motivation for relevant mathematical studies and research over the years [31, 65]. Its challenging theoretical complexity, combined with a wide variety of current problems that require numerical simulations based on this model, justifies such interest. The fundamental model derived in the context of the kinetic theory of gases concerns a nonlinear integro-differential operator. The treatment of the collision integral operator is known to be a significant challenge. For simulation purposes, the Linearized Boltzmann equation (LBE) or kinetic models derived from the original equation are commonly used. On the other hand, when interactions between particles are neglected and the goal is to describe the transport of a particle in a material medium, the balance of the particles results in a linear model. The linear version of the Boltzmann equation is commonly referred to as the Transport Equation (TE) or Radiative Transfer Equation (RTE), for photon applications.
@@
In this work, we report studies and advances related to the solution of the linear Boltzmann equation in one and two spatial dimensions. We discuss relevant aspects of neutron and photon transport modeling, such as in nuclear reactor shielding problems and optical tomography. In addition, we include a brief report on kinetic models of the rarefied gas dynamics, which are relevant, for instance, in the study of micro-electro-mechanical systems, and their connection to the linearized Boltzmann equation (LBE).
@@
The already noted complexity of the Boltzmann models reinforces the development of deterministic methods that can provide high-quality weak-form solutions. The discretization of variables is necessary in both purely numerical approaches and analytical techniques. In this context, the discrete ordinates method is well-known and is used in established codes [43] to deal with angular variables. The solution for selected discrete directions of the particles, along with the numerical approximation of the integral operator by a quadrature scheme, transforms the original problem into a system of first-order partial or ordinary differential equations.
@@
In the literature, the discrete ordinates method is often referred to as the $S_{N}$ method [36]. In fact, in this case, in addition to the angular discretization, spatial variables are also usually discretized. The work of Chandrasekhar [32] in astrophysics was a pioneering effort in proposing analytical solutions for the discrete ordinates approximation of the RTE equation in one-dimensional geometry, expressed in terms of the spatial variable. In more recent years, the Analytical Discrete Ordinates (ADO) Method, [17, 20], a version of Chandrasekhar’s solution,  was proposed. The technique comprises a simultaneous set of features that made Chandrasekhar’s original ideas computationally viable, as well as expanded the applicability of the method to problems in other areas of transport theory, such as neutron transport and rarefied gas dynamics. It is worth mentioning such features: the use of an arbitrary numerical quadrature scheme defined in the half-range, not restricted to the Gauss-Legendre quadrature scheme; the separation constants are obtained through the solution of an eigenvalue problem, instead of searching for roots of polynomials; the resulting eigenvalue problem is of half the order of the number of directions; finally, a scaling introduced to deal with the exponential terms in the solution prevents “overflow”.
@@
In two-dimensional Cartesian geometry, the discrete ordinates approximation reduces the integro-differential equation to a system of first-order partial differential equations. Numerical integration on the unit sphere becomes an issue. In the case of spatial variables, the so-called nodal methods [3] reduce the complexity of the model, following the idea of finite volume schemes, and are known to perform better in coarser meshes. The connection between the spatial and angular grids is usually made in the literature through sweep schemes [43] that demand high computational time.
@@
In this framework, the ADO formulation, along with nodal schemes (ADO-Nodal) method, was proposed for solving the discrete ordinates approximation of the two-dimensional linear Boltzmann (transport) equation [6, 53, 12]. The discrete ordinates equations are transversely integrated over a region (node) of the domain, yielding one-dimensional equations for average angular fluxes or intensities, in $x$ and $y$-directions. The ADO method [17] is then applied to the one-dimensional equations, with approximations for the unknowns (transverse leakage terms) on the contours of the regions, to derive explicit solutions for the spatial variables.
@@
Due to the analytical characteristic of the ADO-Nodal method, as first proposed in [6], the integration procedure to generate the nodal equations was defined for the whole interval of the $x$ and $y$ variables definition, respectively. No domain division into nodes was carried out. Later, the idea of local solutions was considered for the treatment of heterogeneous media problems [12] since different materials constitute the physical domain.
@@
Among the good features of this formulation, which is a spectral technique, one may cite: the solution of the transverse integrated one-dimensional equations, either in $x$ and $y$-directions are explicitly written in terms of the spatial variables along with the fact that the associated eigenvalue problem is of reduced order, to only half of the number of discrete directions. Significant advantages in comparison with the nodal schemes available then. With these features, the methodology applied to a broad class of neutron and radiation transport problems has advantages in providing fast and accurate solutions. Enhanced performance has been noted for coarser meshes if compared with analogous methodologies.
@@
The scheme does not use sweep. Linear systems must be solved to fully establish the general solutions. A significant number of discrete directions in the simulations is theoretically recognized as an error control requirement between the discrete ordinate solution and the exact solution [47]. Furthermore, from a physical perspective, problems in highly anisotropic scattering media, for example, also impose such a condition, which means a choice of high-order quadrature schemes to approximate the integral term of the equation. Quadrature order is a factor that directly affects the linear system. We implemented and analyzed different approaches to solve the linear system associated with the ADO-Nodal formulation. We analyzed the performance of the schemes applied to problems in different areas of particle transport, such as neutrons and radiation, as well as the influence of using different schemes for representing the discrete ordinates [8].
@@
In this work, we present a general overview of the ADO methodology, aiming to highlight advances in the area and contributions to this topic, as already mentioned, of such complexity. We highlight, in particular, different aspects and mathematical tools necessary for the derivation and implementation of the formulation, as well as for error analysis.
@@
The manuscript is organized as follows: the general integro-differential formulation is introduced in Section 2, for photons (RTE). In Sections 3 and 4, the simpler one-dimensional models are discussed and the ADO formulation is derived. Two-dimensional geometry problems are presented in Section 5, where the multidimensional quadrature schemes are commented, and the ADO-Nodal formulation is developed. Next, in Section 6, neutron transport is discussed, highlighting the error in the spatial discretization of the two-dimensional transport equation. Problems in the general area of rarefied gas dynamics, its connection with the LBE, and the application of the ADO method in this area are analyzed in Section 7. A short report on Inverse problems (Section 8) in the field and the final remarks conclude the text.
2 The Integro-differential Formulation
We begin with the time-independent (RTE), for homogeneous media, which results from a balance of gains and losses in the phase space,
@@ 2 The Integro-differential Formulation
Here $I({\bf r},{\bm{\Omega}})$ is the intensity of the radiation in $\bf r$} along the direction ${\bm{\Omega}}=(\mu,\eta,\xi)$ of the particle as a vector in the unit sphere $S$, with $\mu^{2}+\eta^{2}+\xi^{2}=1$, and (see Fig. 1)
@@
where $\Theta$, the phase-angle, is the angle between incident and scattered directions; $C_{l}$ are the expansion coefficients; $L$ refers to the degree of anisotropy of the medium; $P_{l}$ are the Legendre polynomial of lthl-th degree. Using the Addition theorem for the Legendre Polynomials [63]
@@
The Henyey-Greenstein function is widely used to model scattering in biological tissues, as in optical tomography applications. In this case, $C_{l}=(2l+1)g^{l}$, being $g$ the asymmetry factor, $g\in(-1,1)$. That way $g=0$ indicates isotropic scattering while $0<g<1$ indicates forward scattering and 1<g<0-1<g<0 indicates backward scattering. On the other hand, the exact representation of the Henyey-Greenstein function is of the form
@@
To complete the formulation, Eq. (1) is supplemented with boundary conditions. In this work, we introduce them later in the text.
3 A “Simplest” Problem
For an introduction to transport theory, it may be helpful to start with the simplest possible geometry: a one-dimensional homogeneous medium with isotropic scattering. The general coordinate system described in Fig. 1 is now restricted to what we see in Fig. 2. Under such conditions, we consider the problem given by
@@ 3 A “Simplest” Problem
Here $\tau\in\left(\tau_{a},\tau_{b}\right)$, $\tau$ is the optical thickness, $\mu\in\left[-1,1\right]$ is the cosine of the polar angle as measured from the positive $\tau$-axis, and $\varpi$ is the albedo for single scattering. We assume that the incoming intensities, $F_{1}(\mu)$ and $F_{2}(\mu)$ are known on the boundaries. Case’s work [27] was a landmark in transport theory, presenting a rigorous mathematical development for addressing such a problem. This procedure is usually called Case’s method or the singular eigenfunction expansion method. Known as the only exact solution to this problem, despite some additional results [28] in the ’60s and ’70s, mainly, it is only viable in idealized cases. However, this solution has been the basis for the development of numerical and semi-analytical methods. More importantly, it emphasizes the difficulty of solving problems of this type, characterized by singular eigenfunctions, associated with continuous and discrete spectrum, and the need for solutions of singular integral equations in assessing orthogonality conditions.
In other words, to address the fundamental problems in various areas of nuclear science, radiative transfer, and rarefied gas dynamics, alternative approaches needed to be developed. Two approaches stand out: probabilistic, such as the Monte Carlo method [40], and deterministic, which includes the discrete ordinates method.
In the next section, we present the development of the ADO formulation for a still basic problem, with the same geometry as in the previous case, where we introduce complexities such as the treatment of anisotropic media and boundary conditions that include reflection. The idea is to show how much simpler (and powerful at the same time) it is such an approach, in comparison with the Case’s solution.
@@ 4 The ADO Solution for One-dimensional Geometry Problems
In this work, we apply the Analytical Discrete Ordinates (ADO) Method [17] to solve the slab problem defined by Eqs. (7) to (9). Since the ADO formulation was derived in detail for similar models in earlier papers [20, 4, 39], we briefly develop the main steps here. We begin with the equation valid for anisotropic scattering media,
@@
for $\tau\in\left(\tau_{a},\tau_{b}\right)$, $\tau$ is the optical thickness, $\mu\in\left[-1,1\right]$ is the cosine of the polar angle as measured from the positive $\tau$-axis, $\varpi\in[0,1)$ is the albedo for single scattering, $\beta_{l}$ are the coefficients in the $L$th-order expansion of the scattering law and $Q(\tau,\mu)$ is an inhomogeneous source term. We assume boundary conditions defined as
@@
for $\mu\in\left(0,1\right]$, which allows for specular and diffuse reflection, with specular and diffuse reflection coefficients, respectively, $\rho_{i}^{s}$ and $\rho_{i}^{d}$, $i=1,2$. We also assume that there are known incident distributions at the boundaries $\tau=\tau_{a}$ and $\tau=\tau_{b}$, given by $F_{1}\left(\mu\right)$ and $F_{2}\left(\mu\right)$, respectively.
We will express the general solution of Eq. (10) as the sum of the homogeneous plus particular solutions. Following Barichello [4], for developing the homogeneous solution, we assume $Q(\tau,\mu)=0$ and rewrite the integral term in equation (10) as
@@
$\pm\mu_{i}\frac{d}{d\tau}\Phi\left(\tau,\pm\mu_{i}\right)+\Phi\left(\tau,\pm\mu_{i}\right)\\ =\frac{\varpi}{2}\sum_{l=0}^{L}\beta_{l}P_{l}\left(\mu_{i}\right)\sum_{k=1}^{N}w_{k}P_{l}\left(\mu_{k}\right)\left[\Phi\left(\tau,\pm\mu_{k}\right)+\left(-1\right)^{l}\Phi\left(\tau,\mp\mu_{k}\right)\right],$ (14)
@@
Equations (14) represent a first-order ordinary differential system of equations. Thus, we seek exponential solutions of the form
@@
Substituting (15) into equations (14), for $i=1,2,\dots,N$, we obtain
@@
and we add the two sets of equations given in (17), to find that
@@
Since $\mu_{i}>0$ for $i=1,2,\dots,N$, we note that the inverse of the matrix $\mathbf{M}$, denoted by $\mathbf{M}^{-1}$, is well defined. Analogously, we subtract the two sets of equations in (17), to obtain
@@
We note that the separation constants, $\nu$, will appear in pairs $\left(\pm\right)$. In addition, the eigenvalue problems (34) and (35) are of half-order in comparison with the ones obtained in the standard discrete ordinates methods, which are based on full-range quadrature schemes. Continuing, from equation (32) we have
@@
from the solution of the eigenvalue problem, along with the elementary solutions, (39) and (40), we can write the discrete ordinates solution for the homogeneous version of Eq. (10) in the vector form
@@
where $A_{j}$ and $B_{j}$ are the arbitrary coefficients to be determined next. We note that, in writing the solution in this way, we avoid the computational issue of exponential terms which can lead to overflow [20, 4].
To complete our solution, concerning to the particular solutions we follow [10]. We omit here the details of the derivation, which was based on the infinite-medium Green’s function. We express the components of the $N\times 1$ vector
@@
At this point, we can represent, in a vector form, the general solution of Eq. (10),
@@
To completely establish the solution, we need to determine the superposition coefficients, $A_{j}$ and $B_{j}$. These constants can be calculated as the solution of a $2N\times 2N$ linear system of equations, given below, obtained from the application of the boundary conditions, Eqs. (11) and (12),